# KiDS-Legacy: Consistency of cosmic shear measurements and joint cosmological constraints with external probes

### v1

@@ 4.2 Splits at the catalogue level
Figure 3: COSEBIs E-mode measurements and their best-fitting model for a split cosmological analysis of the North-South split catalogue. The green and purple data points show the measurements of the KiDS-North and KiDS-South sample, respectively. The best-fitting theoretical predictions are given by the solid lines, and the $1\sigma$ interval of the translated posterior distributions are illustrated by the shaded regions. Each panel represents auto- or cross-correlation between tomographic bins. For visualisation purposes, we display the discrete $n$ modes with an offset on the x-axis.We note that the E-mode signals are highly correlated within a tomographic bin and advise against a so-called ?$\chi$-by-eye?.
In this section, we split the data on the catalogue level, analysing the two distinct sections of the KiDS footprint, KiDS-North and KiDS-South, and splitting the galaxy sample by colour. When conducting the likelihood analysis, we construct a joint data vector and covariance matrix of both subsets, doubling the dimensionality of the data vector with respect to the fiducial analysis. In Appendix A we report the data properties of each catalogue split with calibrated redshifts and shear measurements for each sample along with the B-mode signal of each subset which we find to be consistent with the null hypothesis.
@@ 4.2.1 North-South split
Table 3: Consistency metrics for catalogue level splits of KiDS-Legacy data.
@@
Red vs Blue $T_{\rm B}=3.0$0.44 2.61 1.08 0.08 0.74 0.76 0.40
Red vs Blue $T_{\rm B}=1.9$1.29 2.84 0.70 0.46 0.25 0.64 0.35
 [^8] The second and third columns report the Bayes ratio and the tension level inferred from the suspiciousness, respectively. The next three columns present the results of the tier 2 test for $\Omega_{\rm m}$, $S_{8}$, and their combination. The final two columns shows the tension level arising from the tier 3 test in terms of the $p$-value for data vector predictions for the first subset of the catalogue inferred from the TPD of the other subset (column 7) and vice versa (column 8).
@@
As discussed in Sect. 2, the KiDS observations were taken on two distinct patches on sky: one at the celestial equator, dubbed KiDS-North, and one at the South Galactic Pole, dubbed KiDS-South. The two patches, which are shown in fig. 2 of W25, cover a similar area on sky with approximately $496\,{\rm deg}^{2}$ of post-masking data in KiDS-North and $472\,{\rm deg}^{2}$ in KiDS-South. Thus, both patches contain a comparable number of sources per tomographic bin. In the fiducial analysis pipeline, we combine independent measurements of the cosmic shear 2PCF per patch into a single measurement, from which we compute the COSEBIs data vector. In the North-South split, we keep the 2PCF measurements in each patch separate in order to test their consistency. By doing so, we obtain one data vector per hemisphere. Since the two patches are separated on sky, we do not expect a cross-correlation signal between patches. Therefore, the combined covariance matrix only consists of two non-zero blocks, each containing the covariance for North and South, respectively.
@@
In Fig. 3, we present the COSEBIs data vector per hemisphere along with the corresponding theory prediction from the best-fitting model with uncertainties inferred from the TPDs. As the data properties of the KiDS-North and KiDS-South patches are very similar, in terms of the redshift distributions, multiplicative shear and redshift calibration, and the ellipticity dispersion (see Appendix A), we have chosen to use a single shared set of observational nuisance parameters in our likelihood analysis, shown in the left panel of Fig. 4. Although the KiDS-North patch tends to favour a shift towards lower values of $\Omega_{\rm m}$ compared to our fiducial analysis, both patches are in good agreement with a value of $N_{\sigma}(\Delta\Omega_{\rm m})=1.09$ inferred in the tier 2 multi-parameter consistency test. Considering $S_{8}$, we find both patches to be in agreement, which is confirmed by the tier 1 evidence and tier 3 PPD tests tabulated in the first row in Table 3. We conclude that the cosmological constraints from observations in KiDS-North and KiDS-South are fully consistent.
@@
Figure 4: Posterior distribution of parameter duplicates in the $\Omega_{\rm m}-S_{8}$ plane for catalogue level splits for COSEBIs. Left panel: North-South split. Middle panel: Red-blue split defined via a cut on the spectral type of $T_{\rm B}=3.0$. Right panel: Red-blue split defined via a cut on the spectral type of $T_{\rm B}=1.9$. For reference, the black dashed contours show constraints from the analysis with a single set of parameters modelling both data subsets. The inner and outer contours of the marginalised posteriors correspond to the 68% and 95% credible intervals, respectively.
@@ 4.2.2 Red-blue split
Figure 5: COSEBIs E-mode measurements and their best-fitting model for a split cosmological analysis of the red-blue split catalogue, defined via a cut on the spectral type $T_{\rm B}=3.0$. The red and blue data points show the measurements of the red and blue sample, respectively. The best-fitting theoretical predictions are given by the solid lines, and the $1\sigma$ interval of the translated posterior distributions are illustrated by the shaded regions. Each panel represents auto- or cross-correlation between tomographic bins, as indicated by the label in the top right corner. For visualisation purposes, we display the discrete $n$ modes with an offset on the x-axis. We note that the E-mode signals are highly correlated within a tomographic bin and advise against a so-called ?$\chi$-by-eye?.
@@
We follow the methodology of Li et al. (2021), dividing the KiDS-Legacy sample into two subsets based on the spectral type $T_{\rm B}$ reported by the bpz code. Li et al. (2021) defines blue galaxies via a threshold of $T_{\rm B}>3.0$ to provide a similar constraining power per data subset. This is in contrast to appendix B of W25, where red galaxies are selected with a threshold of $T_{\rm B}\leq 1.9$. We present an analysis of each threshold [^9] We note another alternative, that we do not explore here, is to split a catalogue by colour as proposed by McCullough et al. (2024), applying a SOM-based selection on $r-z$ colour in order to derive a high-purity sample of blue galaxies., where dividing the sample at a threshold of $T_{\rm B}=3.0$ results in the red sample containing approximately one third of the galaxies, with the more stringent cut of $T_{\rm B}=1.9$ leaving $\sim 16\%$ of galaxies in the red sample. In Appendix A, we present the redshift distributions of each sample and the separate redshift nuisance parameters that are marginalised over in our analysis. In contrast to the division of galaxies by hemisphere, the red and blue galaxy samples are expected to be highly correlated and these correlations are taken into account via the cross-covariance between cosmic shear measurements of the red and blue samples computed with the OneCovariance code.
@@
Our fiducial cosmic shear analysis adopts the NLA-$M$ IA model which sets any blue galaxy alignment to zero for $T_{\rm B}>1.9$. This model is therefore not applicable when considering a split between red and blue galaxies at a threshold of $T_{\rm B}=3.0$ and we revert to the standard NLA model in this analysis for both $T_{\rm B}$ thresholds with a separate amplitude parameter $A_{\rm IA}$ for each galaxy sample.
@@
The measured COSEBIs E modes for the colour-based split with a threshold $T_{\rm B}=3.0$ and the TPDs from the best-fitting theory model are illustrated in Fig. 5. Here, each colour represents measurements of the auto- and cross-correlation between tomographic bins of the given sample. We chose not to include cross-correlation measurements between the red and blue bins as these signals mix contributions from theoretical predictions for the red and blue signals which can not be easily modelled with our current pipeline. The cross-correlation between the red and blue data points is, however, taken into account in our consistency analysis via the joint covariance matrix.
@@
The posterior distribution for both instances of $\Omega_{\rm m}$ and $S_{8}$ in the split cosmological analysis are shown in the middle and right panels of Fig. 4. As expected, the red galaxy sample yields weaker constraints on $S_{8}$ and $\Omega_{\rm m}$ than the blue sample due to the higher number density of blue galaxies. This is particularly true for the $T_{\rm B}=1.9$ selected sample where the marginalised posteriors are almost unconstrained by the red galaxy sample. Nevertheless, the cosmological parameter posterior distributions for both red/blue splits are in good agreement. Looking at the consistency metrics listed in Table 3 for the $T_{\rm B}=3.0$ ($T_{\rm B}=1.9$) selected samples, we find that the suspiciousness test shows a preference for the split cosmological model at $2.61\sigma$ ($2.84\sigma$). In contrast, the tier 2 test yields an agreement between parameters with $N_{\sigma}\leq 1.08$ ($N_{\sigma}\leq 0.70$) when considering $S_{8}$, $\Omega_{\rm m}$, and their combination. This is confirmed by the tier 3 tests, which suggests a good agreement between the TPDs and the observed data in both subsets.
@@
The tier 1 evidence-based preference for the split cosmological model can be explained by the additional freedom in the modelling of intrinsic galaxy alignments. While the combined galaxy sample analysis assumes a shared IA amplitude for both red and blue galaxies, the split analysis models intrinsic alignments with two independent parameters. As shown in Fig. 6, there is a significant difference between the IA amplitudes for the two samples with a marginal mode and the 1D 68% highest posterior density interval (HPDI) of $A^{\rm blue}_{\rm IA}=-0.75^{+0.43}_{-0.62}$ and $A^{\rm red}_{\rm IA}=1.08^{+0.46}_{-0.44}$ for the $T_{\rm B}=3.0$ sample split. For the $T_{\rm B}=1.9$ split, we find the IA amplitude of blue galaxies to be $A^{\rm blue}_{\rm IA}=-0.32^{+0.33}_{-0.36}$, while the red galaxy sample yields $A^{\rm red}_{\rm IA}=3.32^{+1.13}_{-0.97}$. This corresponds to a difference in their posterior distribution at $N_{\sigma}(\Delta A_{\rm IA})=2.81$ for $T_{\rm B}=3.0$ and $N_{\sigma}(\Delta A_{\rm IA})=2.57$ for $T_{\rm B}=1.9$. These results are compatible with the central assumption of the NLA-$M$ model used in our fiducial cosmic shear analysis, which assumes zero alignment of blue galaxies. We therefore do not interpret the tier 1 result as an indication of internal inconsistency given that the physical mechanism behind red and blue galaxy alignment is expected to differ. As the evidence-based tier 1 consistency test compresses the full posterior into a single statistic, the physical difference between red and blue galaxies is reflected as a preference for the split cosmological model. Based on the other consistency metrics reported in Table 3, we can therefore conclude there is internal consistency between our samples of red and blue galaxies.
@@ 4.3 Consistency between summary statistics
While the two primary summary statistics in the fiducial analysis, COSEBIs and band powers, as well as the additional statistic, binned 2PCFs, are derived from the same cosmic shear 2PCF measurements, they differ in their sensitivity to spatial scales as well as systematic and modelling effects. Thus, we quantify the consistency between all combinations of two summary statistics. Although this does not correspond to a split of the data vector or the catalogue since each statistic originates from the same measurements of the cosmic shear 2PCFs, we can nevertheless use the same methodology to quantify whether or not the statistics prefer different cosmologies. In practice, we construct three data vectors that each combine measurements of two summary statistics and model the theoretical prediction with one set of parameters per statistic. As depicted in Eq. (1), they differ only in the corresponding weight function, which vary in their sensitivity to different angular scales. Therefore, we expect the statistics to be highly correlated, since the different summary statistics are derived from the same two-point correlation function measurements. Thus, it is particularly important to derive a robust estimate of the covariance between summary statistics, which is enabled by the OneCovariance code. The resulting correlation matrix is displayed in Fig. 7. As discussed in Reischke et al. (2024), the full covariance matrix between all summary statistics can be non-positive definite due to numerical noise. However, the sub-covariance matrices between two summary statistics are still positive definite and invertible.
@@
The marginalised posterior distribution for $\Omega_{\rm m}$ and $S_{8}$ are displayed in Fig. 8. Here, the solid contours refer to the analysis with the split cosmological model and the black dashed lines show constraints from the analysis of the combined data vector of two summary statistics with a single set of parameters. For the combination of COSEBIs and 2PCFs as well as band powers and 2PCFs, we find the constraints from the combined analysis to be in agreement with the split analyses. However, for the combination of COSEBIs and band powers we observe a preference for lower values of $\Omega_{\rm m}$ in the combined analysis. We further investigate the origin of this feature by decomposing the likelihood into the contribution from the auto-correlation of COSEBIs and band powers and the contribution from their cross-correlation. This analysis shows that the shift towards low $\Omega_{\rm m}$ is driven by the cross-covariance terms between COSEBIs and band powers. Additionally, we inspect the posterior distribution of the remaining split parameters, which is displayed in Fig. 9. We find that the two instances of the baryon feedback parameter, $\log T_{\rm AGN}$, show a preference for different amounts of baryonic feedback. While the COSEBIs posterior peaks at a low value of $\log T_{\rm AGN}$, corresponding to a dark matter only scenario, the posterior from band powers tends towards the upper edge of the prior. The origin of this feature most likely lies in their varying response to different scales. In the combined analysis with a single set of parameters, in which both statistics share the same baryonic feedback parameter, this most likely causes the shift of the posterior towards low $\Omega_{\rm m}$. As a consequence, the evidence-based consistency metric, provided in Table 4, reports a preference for the split cosmological model with $N_{\sigma,S}=3.78$. In the tier 2 parameter space and tier 3 data space metrics, however, we find both statistics to be in agreement. We note that our default consistency test in parameter space focusses on the two parameters that are mostly constrained by our cosmic shear data, $\Omega_{\rm m}$ and $S_{8}$. Considering the apparent discrepancy in the baryon feedback parameter, we compute the significance of the shift in $\log T_{\rm AGN}$, which we find to be $N_{\sigma}(\Delta\log T_{\rm AGN})=2.54$. Furthermore, we note that the posterior in the combined analysis of COSEBIs and band powers closely resembles the posterior of an analysis with 2PCFs (see appendix E in W25). As can be inferred from the corresponding window functions, the combination of COSEBIs and band powers covers approximately the same range of scales that is probed by 2PCFs, which provides an explanation for the similarity between their posteriors.
@@
For the remaining combinations of summary statistics, our consistency analysis finds an agreement in all tests. Therefore, we conclude that for the parameters of interest we find the three summary statistics to be in agreement. This confirms the result of the fiducial cosmic shear analysis (W25), which reports a good agreement between marginalised $S_{8}$ constraints inferred individually with the three statistics. However, we note that a combined analysis of two summary statistics proves to be challenging due to the high degree of correlation between the statistics, as showcased in our combined analysis of COSEBIs and band powers.
Table 4: Consistency metrics for the combination of summary statistics.
@@
 [^10] The second and third columns report the Bayes ratio and the tension level inferred from the suspiciousness, respectively. The next three columns present the results of the tier 2 test for $\Omega_{\rm m}$, $S_{8}$, and their combination. The final two columns shows the tension level arising from the tier 3 test in terms of the $p$-value for data vector predictions for the first statistic inferred from the TPD of the other statistic (column 7) and vice versa (column 8).

### v2

@@ 4.2 Splits at the catalogue level
Figure 3: COSEBIs E-mode measurements and their best-fitting model for a split cosmological analysis of the North-South split catalogue. The green and purple data points show the measurements of the KiDS-North and KiDS-South sample, respectively. The best-fitting theoretical predictions are given by the solid lines, and the $1\sigma$ interval of the TPDs are illustrated by the shaded regions. Each panel represents auto- or cross-correlation between tomographic bins. For visualisation purposes, we display the discrete $n$ modes with an offset on the x-axis.We note that the E-mode signals are highly correlated within a tomographic bin and advise against a so-called $\chi$-by-eye?.
In this section, we split the data on the catalogue level, analysing the two distinct sections of the KiDS footprint, KiDS-North, and KiDS-South, and splitting the galaxy sample by colour. When conducting the likelihood analysis, we construct a joint data vector and covariance matrix of both subsets, doubling the dimensionality of the data vector with respect to the fiducial analysis. In Appendix A, we report the data properties of each catalogue split with calibrated redshifts and shear measurements for each sample along with the B-mode signal of each subset, which we found to be consistent with the null hypothesis.
@@ 4.2.1 North-South split
Table 3: Consistency metrics for catalogue-level splits of KiDS-Legacy data. [^8] The second and third columns report the Bayes ratio and the tension level inferred from the suspiciousness, respectively. The next three columns present the results of the tier 2 test for $\Omega_{\rm m}$, $S_{8}$, and their combination. The final two columns shows the tension level arising from the tier 3 test in terms of the $p$-value for data vector predictions for the first subset of the catalogue inferred from the TPD of the other subset (column 7) and vice versa (column 8).
@@
Red vs blue, $T_{\rm B}=3.0$ -0.44 2.61 1.08 0.08 0.74 0.76 0.40
Red vs blue $T_{\rm B}=1.9$ -1.29 2.84 0.70 0.46 0.25 0.64 0.35
@@
As discussed in Sect. 2, the KiDS observations were taken on two distinct patches on sky: one at the celestial equator, dubbed KiDS-North, and one at the South Galactic Pole, dubbed KiDS-South. The two patches, which are shown in fig. 2 in W25, cover a similar area on sky with approximately $496\,{\rm deg}^{2}$ of post-masking data in KiDS-North and $472\,{\rm deg}^{2}$ in KiDS-South. As a result, both patches contain a comparable number of sources per tomographic bin. In the fiducial analysis pipeline, we combined independent measurements of the cosmic shear 2PCF per patch into a single measurement, from which we computed the COSEBIs data vector. In the North-South split, we kept the 2PCF measurements in each patch separate in order to test their consistency. By doing so, we obtained one data vector per hemisphere. Since the two patches are separated on sky, we did not expect a cross-correlation signal between patches. Therefore, the combined covariance matrix only consists of two non-zero blocks, each containing the covariance for North and South, respectively.
@@
In Fig. 3, we present the COSEBIs data vector per hemisphere along with the corresponding theory prediction from the best-fitting model with uncertainties inferred from the TPDs. As the data properties of the KiDS-North and KiDS-South patches are very similar, in terms of the redshift distributions, multiplicative shear and redshift calibration, and the ellipticity dispersion (see Appendix A), we have chosen to use a single shared set of observational nuisance parameters in our likelihood analysis, shown in the left panel of Fig. 4. Although the KiDS-North patch tends to favour a shift towards lower values of $\Omega_{\rm m}$ compared to our fiducial analysis, both patches are in good agreement with a value of $N_{\sigma}(\Delta\Omega_{\rm m})=1.09$ inferred in the tier 2 multi-parameter consistency test. Considering $S_{8}$, we found both patches to be in agreement, which is confirmed by the tier 1 evidence and tier 3 PPD tests tabulated in the first row in Table 3. We conclude that the cosmological constraints from observations in KiDS-North and KiDS-South are fully consistent.
@@
Figure 4: Posterior distribution of parameter duplicates in the $\Omega_{\rm m}-S_{8}$ plane for catalogue-level splits for COSEBIs. Left panel: North-South split. Middle panel: Red-blue split defined via a cut on the spectral type of $T_{\rm B}=3.0$. Right panel: Red-blue split defined via a cut on the spectral type of $T_{\rm B}=1.9$. For reference, the black dashed contours show constraints from the analysis with a single set of parameters modelling both data subsets. The inner and outer contours of the marginalised posteriors correspond to the 68% and 95% credible intervals, respectively.
@@ 4.2.2 Red-blue split
Figure 5: COSEBIs E-mode measurements and their best-fitting model for a split cosmological analysis of the red-blue split catalogue, defined via a cut on the spectral type $T_{\rm B}=3.0$. The red and blue data points show the measurements of the red and blue sample, respectively. The best-fitting theoretical predictions are given by the solid lines, and the $1\sigma$ interval of the TPDs are illustrated by the shaded regions. Each panel represents auto- or cross-correlation between tomographic bins, as indicated by the label in the top right corner. For visualisation purposes, we display the discrete $n$ modes with an offset on the x-axis. We note that the E-mode signals are highly correlated within a tomographic bin and advise against a so-called $\chi$-by-eye?.
@@
We followed the methodology of Li et al. (2021), dividing the KiDS-Legacy sample into two subsets based on the spectral type $T_{\rm B}$ reported by the bpz code. Li et al. (2021) defined blue galaxies via a threshold of $T_{\rm B}>3.0$ to provide a similar constraining power per data subset. This is in contrast to appendix B in W25, where red galaxies are selected with a threshold of $T_{\rm B}\leq 1.9$. We present an analysis of each threshold [^9] We note another alternative, that we do not explore here, is to split a catalogue by colour as proposed by McCullough et al. (2024), applying a SOM-based selection on rzr-z colour in order to derive a high-purity sample of blue galaxies., where dividing the sample at a threshold of $T_{\rm B}=3.0$ results in the red sample containing approximately one third of the galaxies, with the more stringent cut of $T_{\rm B}=1.9$ leaving $\sim 16\%$ of galaxies in the red sample. In Appendix A, we present the redshift distributions of each sample and the separate redshift nuisance parameters that are marginalised over in our analysis. In contrast to the division of galaxies by hemisphere, the red and blue galaxy samples are expected to be highly correlated and these correlations are taken into account via the cross-covariance between cosmic shear measurements of the red and blue samples computed with the OneCovariance code.
@@
Our fiducial cosmic shear analysis adopts the NLA-$M$ IA model, which sets any blue galaxy alignment to zero for $T_{\rm B}>1.9$. Thus, this model is not applicable when considering a split between red and blue galaxies at a threshold of $T_{\rm B}=3.0$, which requires a recalibration of the IA mass scaling. Since this is beyond the scope of this work, we revert to the standard NLA model in this analysis for both $T_{\rm B}$ thresholds with a separate amplitude parameter, $A_{\rm IA}$, for each galaxy sample. We note that as discussed in W25, the cosmological constraints are highly consistent between different IA models. Therefore, we do not expect the choice of IA model to make an impact on the internal consistency test in cosmological parameter space.
@@
The measured COSEBIs E modes for the colour-based split with a threshold $T_{\rm B}=3.0$ and the TPDs from the best-fitting theory model are illustrated in Fig. 5. Here, each colour represents measurements of the auto- and cross-correlation between tomographic bins of the given sample. We chose not to include cross-correlation measurements between the red and blue bins as these signals mix contributions from theoretical predictions for the red and blue signals that cannot be easily modelled with our current pipeline. The cross-correlation between the red and blue data points is, however, taken into account in our consistency analysis via the joint covariance matrix.
@@
The posterior distribution for both instances of $\Omega_{\rm m}$ and $S_{8}$ in the split cosmological analysis are shown in the middle and right panels of Fig. 4. As expected, the red galaxy sample yields weaker constraints on $S_{8}$ and $\Omega_{\rm m}$ than the blue sample due to the higher number density of blue galaxies. This is particularly true for the $T_{\rm B}=1.9$ selected sample where the marginalised posteriors are almost unconstrained by the red galaxy sample. Nevertheless, the cosmological parameter posterior distributions for both red-blue splits are in good agreement. Looking at the consistency metrics listed in Table 3 for the $T_{\rm B}=3.0$ ($T_{\rm B}=1.9$) selected samples, we found that the suspiciousness test shows a preference for the split cosmological model at $2.61\sigma$ ($2.84\sigma$). In contrast, the tier 2 test yields an agreement between parameters with $N_{\sigma}\leq 1.08$ ($N_{\sigma}\leq 0.70$) when considering $S_{8}$, $\Omega_{\rm m}$, and their combination. This is confirmed by the tier 3 tests, which suggests a good agreement between the TPDs and the observed data in both subsets.
@@
The tier 1 evidence-based preference for the split cosmological model can be explained by the additional freedom in the modelling of intrinsic galaxy alignments. While the combined galaxy sample analysis assumes a shared IA amplitude for both red and blue galaxies, the split analysis models intrinsic alignments with two independent parameters. As shown in Fig. 6, there is a significant difference between the IA amplitudes for the two samples with a marginal mode and the 1D 68% highest posterior density interval of $A^{\rm blue}_{\rm IA}=-0.75^{+0.43}_{-0.62}$ and $A^{\rm red}_{\rm IA}=1.08^{+0.46}_{-0.44}$ for the $T_{\rm B}=3.0$ sample split. For the $T_{\rm B}=1.9$ split, we found the IA amplitude of blue galaxies to be $A^{\rm blue}_{\rm IA}=-0.32^{+0.33}_{-0.36}$, while the red galaxy sample yields $A^{\rm red}_{\rm IA}=3.32^{+1.13}_{-0.97}$. This corresponds to a difference in their posterior distribution at $N_{\sigma}(\Delta A_{\rm IA})=2.81$ for $T_{\rm B}=3.0$ and $N_{\sigma}(\Delta A_{\rm IA})=2.57$ for $T_{\rm B}=1.9$. These results are compatible with the central assumption of the NLA-$M$ model used in our fiducial cosmic shear analysis, which assumes zero alignment of blue galaxies. We therefore do not interpret the tier 1 result as an indication of internal inconsistency given that the physical mechanism behind red and blue galaxy alignment is expected to differ. As the evidence-based tier 1 consistency test compresses the full posterior into a single statistic, the physical difference between red and blue galaxies is reflected as a preference for the split cosmological model. Based on the other consistency metrics reported in Table 3, we can therefore conclude there is internal consistency between our samples of red and blue galaxies.
@@ 4.3 Consistency between summary statistics
While the two primary summary statistics in the fiducial analysis, COSEBIs and band powers, as well as the additional statistic, binned 2PCFs, are derived from the same cosmic shear 2PCF measurements, they differ in their sensitivity to spatial scales as well as systematic and modelling effects. Thus, we quantified the consistency between all combinations of two summary statistics. Although this does not correspond to a split of the data vector or the catalogue since each statistic originates from the same measurements of the cosmic shear 2PCFs, we can nevertheless use the same methodology to quantify whether or not the statistics prefer different cosmologies. In practice, we construct three data vectors that each combine measurements of two summary statistics and model the theoretical prediction with one set of parameters per statistic. As depicted in Eq. (1), they differ only in the corresponding weight function, which vary in their sensitivity to different angular scales. Therefore, we expect the statistics to be highly correlated, since the different summary statistics are derived from the same two-point correlation function measurements. Thus, it is particularly important to derive a robust estimate of the covariance between summary statistics, which is enabled by the OneCovariance code. The resulting correlation matrix is displayed in Fig. 7. As discussed in Reischke et al. (2025), the full covariance matrix between all summary statistics can be non-positive definite due to numerical noise. However, the sub-covariance matrices between two summary statistics are still positive definite and invertible.
@@
The marginalised posterior distributions for $\Omega_{\rm m}$ and $S_{8}$ are displayed in Fig. 8. Here, the solid contours refer to the analysis with the split cosmological model and the black dashed lines show constraints from the analysis of the combined data vector of two summary statistics with a single set of parameters. For the combination of COSEBIs and 2PCFs as well as band powers and 2PCFs, we found the constraints from the combined analysis to be in agreement with the split analyses. However, for the combination of COSEBIs and band powers we observe a preference for lower values of $\Omega_{\rm m}$ in the combined analysis. We further investigated the origin of this feature by decomposing the likelihood into the contribution from the auto-correlation of COSEBIs and band powers and the contribution from their cross-correlation. This analysis shows that the shift towards low $\Omega_{\rm m}$ is driven by the cross-covariance terms between COSEBIs and band powers. Additionally, we inspected the posterior distribution of the remaining split parameters, displayed in Fig. 9. We found that the two instances of the baryon feedback parameter, $\log T_{\rm AGN}$, show a preference for different amounts of baryonic feedback. While the COSEBIs posterior peaks at a low value of $\log T_{\rm AGN}$, corresponding to a dark matter only scenario, the posterior from band powers tends towards the upper edge of the prior. The origin of this feature most likely lies in their varying response to different scales. In the combined analysis with a single set of parameters, in which both statistics share the same baryonic feedback parameter, this most likely causes the shift of the posterior towards low $\Omega_{\rm m}$. As a consequence, the evidence-based consistency metric, provided in Table 4, reports a preference for the split cosmological model with $N_{\sigma,S}=3.78$. In the tier 2 parameter space and tier 3 data space metrics, however, we find both statistics to be in agreement. We note that our default consistency test in parameter space focusses on the two parameters that are mostly constrained by our cosmic shear data, $\Omega_{\rm m}$ and $S_{8}$. Considering the apparent discrepancy in the baryon feedback parameter, we computed the significance of the shift in $\log T_{\rm AGN}$, which we found to be $N_{\sigma}(\Delta\log T_{\rm AGN})=2.54$. Furthermore, we note that the posterior in the combined analysis of COSEBIs and band powers closely resembles the posterior of an analysis with 2PCFs (see appendix F in W25). As can be inferred from the corresponding window functions, the combination of COSEBIs and band powers covers approximately the same range of scales that is probed by 2PCFs, which provides an explanation for the similarity between their posteriors.
@@
For the remaining combinations of summary statistics, our consistency analysis finds an agreement in all tests. Therefore, we conclude that for the parameters of interest, we find the three summary statistics to be in agreement. This confirms the result of the fiducial cosmic shear analysis (W25), which reports a good agreement between marginalised $S_{8}$ constraints inferred individually with the three statistics. However, we note that a combined analysis of two summary statistics proves to be challenging due to the high degree of correlation between the statistics, as showcased in our combined analysis of COSEBIs and band powers.
Table 4: Consistency metrics for the combination of summary statistics. [^10] The second and third columns report the Bayes ratio and the tension level inferred from the suspiciousness, respectively. The next three columns present the results of the tier 2 test for $\Omega_{\rm m}$, $S_{8}$, and their combination. The final two columns shows the tension level arising from the tier 3 test in terms of the $p$-value for data vector predictions for the first statistic inferred from the TPD of the other statistic (column 7) and vice versa (column 8).